8. Construct the marginal product schedule from the following production function.

LABOR

TOTAL PRODUCT

1

3

2

8

3

15

4

21

5

25

6

26

7

24
Answer #1

Marginal Product schedule can be developed as under:

Labor, L Total Product, Q Marginal Product of labor, MPL=Change in Q/Change in L
1 3 3
2 8 (8-3)/(2-1)=5
3 15 (15-8)/(3-2)=7
4 21 (21-15)/(4-3)=6
5 25 (25-21)/(5-4)=4
6 26 (26-25)/(6-5)=1
7 24 (24026)/(7-6)=-2